250 is invested at a discount rate of 18%p.a convertible monthly for the first 3 months followed by an interest rate of 20%p.a convertible quarterly for the first 9 months. Calculate the accumulated sum at the end of the year


1
Expert's answer
2022-01-24T13:35:05-0500

"pv=fv(\frac{1}{(1+\frac{r}{m})^{nm}})\\
250=fv(\frac{1}{(1+\frac{0.18}{12})^{3\times12}})\\
250=fv(\frac{1}{(1.015)^{36}})\\
\frac{250}{0.0.5851}=\frac{0.5851fv}{0.5851}\\
fv=427.28\\
then,\\
427.28=fv(\frac{1}{(1+\frac{0.2}{12})^{\frac{9}{3}}})\\
\frac{427.28}{0.9362}=\frac{0.9362fv}{0.9362}\\
fv=456.40"